Circular No. 8560

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 N2 (HERGENROTHER) D. Herald, Kambah (near Canberra, A.C.T.), reports his recovery of

comet P/1998 W2 on CCD images obtained with a 0.36-m f/3.9 Schmidt-Cassegrain reflector on July 4 and 5 (with astrometry, the following orbital elements, and an ephemeris published on MPEC 2005-N28). Herald notes that stacking fifteen 5-min frames shows a tail of length ≈ 20′′ in p.a. 270^{o}. The indicated correction to the predictions on MPC 45658 and in the 2005 Comet Handbook is ΔT = +0.27 day.

Epoch = 2005 Nov. 6.0 TT

T = 2005 Nov. 2.4811 TT  ω = 13.^{o}8784
e = 0.607511  Ω = 356.5030 }2000.0
q = 1.425773 AU  i = 21.8941

a = 3.632648 AU n^{o} = 0.1423538 P = 6.924 years

COMET C/2005 N3 (LARSON) S. Larson reports his discovery of a comet (discovery observation given

below) on CCD images taken in the course of the Mt. Lemmon Survey, the object showing a 4′′ coma and a 10′′ tail in p.a. 230^{o} on July 5.4 UT. R. H. McNaught obtained confirming CCD images with the 1.0-m f/8 reflector at Siding Spring on July 5.6 that show a condensation merging into a tail 15′′ long in p.a. 240^{o}.

2005 UT  α_{2000}  δ_{2000}  Mag.
July 5.38079  21^{h}23^{m}26.^{s}08  −14^{o}08′44′′.7  20.2

The astrometry (including prediscovery observations from Mount Lemmon on July 3.4), the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2005-N29. The comet is quite likely to be of short period.

T = 2005 Dec. 18.860 TT  ω = 98.^{o}198
Ω = 300.859 }2000.0
q = 1.25088 AU  i = 8.243

COMET 21P/GIACOBINI-ZINNER Total visual magnitude estimates by J. J. Gonzalez (Leon, Spain, 0.20-

m refl.): May 18.12 UT, 11.2; June 2.09, 10.5; 6.11, 10.2; July 3.11, 9.4.

2005 July 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green